OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (-6, 3). y = 3/2x - 1 y = 3/2x + 3 y = -2/3x - 1 y = -2/3x + 3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, perpendicular lines have OPPOSITE slope. So what two options can we eliminate?

OpenStudy (anonymous):

c and d?

OpenStudy (igreen):

No, that's if we were looking for parallel lines, C and D are the one's we want because they have opposite slope to y = 3/2x + 4..

OpenStudy (anonymous):

ohhh ok!

OpenStudy (igreen):

Okay, so we have to plug in (-6, 3) into both of the equations to see which one it is: y = -2/3x - 1 3 = -2/3(-6) - 1 Multiply: 3 = 4 - 1 Can you subtract that?

OpenStudy (anonymous):

3

OpenStudy (igreen):

Yep, so you get 3 = 3. Which means y = -2/3x - 1 is the correct answer.

OpenStudy (anonymous):

thank you!